I got this wrong, can someone help walk me through how these are done. I hate word problems. Missed 6 of them.
Combined population of City A and City b was 25,000 in 1990. By 1995, the population of city a increased 10% and city b increased 9%. Total population increased by 2380 people, so what was the population in 1990?

Answer by ankor@dixie-net.com(12701)

Combined population of City A and City b was 25,000 in 1990. By 1995, the population of city a increased 10% and city b increased 9%. Total population increased by 2380 people, so what was the population in 1990?
:
Let x = population of City A
Then
(25000-x) = population of City B
:
A simple equation:
.10x + .09(25000-x) = 2380
:
.1x + 2250 - .09x = 2380
:
.1x - .09x = 2380 - 2250
:
.01x = 130
:
x = 130/.01
:
x = 13000 is the population of city A in 1990
:
25000 - 13000 = 12000 is the population of city B in 1990
:
:
Check solutions:
.1(13000) + .09(12000) =
1300 + 1080 = 2380 as given